40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 26 22:05:06 2020, gautam, Update, IOO, IMC checks 
    Reply  Wed Feb 26 23:50:51 2020, gautam, Update, IOO, IMC checks sliderCal.pdfAO_inputTFs.pdf
Message ID: 15229     Entry time: Wed Feb 26 23:50:51 2020     In reply to: 15227
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC checks 

In the style of the KA characterization of the CM board, the AO path gain EPICS slider (IN2) of the IMC servo board was stepped by 1 dB through the full available range of -32 dB to +31 dB. For each value of the requested gain, I measured the TF from the injected signal (to IN2) to TP1A on the IMC servo board. I used the BNC connector for this test, whereas we use the LEMO connector for the AO path. The source was tee-d off at the SR785 side, with one leg going to IN2 of the IMC servo board, and the other going to CH1A of the SR785. TP1A of the IMC board was connected to CH2A of the SR785. 

Attachment #1 - Measured gain vs requested gain.

  • When debugging the CM board, it was this kind of test that revealed the faulty latch ICs.
  • -12 dB to -11 dB gain step looks anomalous, but overall the trend seems linear.
  • I was confused by why there should be a discontinuity at this stage of the gain stepping - seems like the scanning script I use changes the SR785 excitation amplitude at this point (from 300mV to 100mV). But why should the size of the excitation signal change the magnitude of the transfer function? Is this indicative of some loading issue?
  • There is an overall offset between the requested gain and measured gain of ~2-3 dB. This seems large.
  • There is nothing in the schematic which would have me expect this - there is a 1/2 divider at the positive input of the differential receiving stage, but this just cancels out the non-inverting gain of x2.

Attachment #2 - Frequency dependent transfer functions

  • There seem to be two families of curves - they correspond to <-12 dB and >-12 dB.
  • The feature at 90 kHz is strange - need to look at the schematic to see what this could be.

The motivation here is to try and figure out why I cannot engage the AO path smoothly in the CARM handoff part of lock acquisiton. I plan to use this information to do some loop modeling and project laser frequency noise coupling in various stages of the lock acquisition process.

Attachment 1: sliderCal.pdf  110 kB  Uploaded Thu Feb 27 09:49:33 2020  | Show | Show all
Attachment 2: AO_inputTFs.pdf  899 kB  Uploaded Thu Feb 27 09:49:40 2020  | Show | Show all
ELOG V3.1.3-